OpenStudy (anonymous):

Which strategy best explains how to solve this problem? Claire bought a cup of coffee and two magazines each Saturday. The coffee cost $2.50 and the magazines cost $3.99 each. What is the total amount of money that Claire had spent after 7 Saturdays?

OpenStudy (anonymous):

A. Work backward. Start with the price of a two magazines ($7.98). Add the cost of a cup of coffee and then divide by 7. B. Make a table. Write the numbers 1 to 7 in the top row of the table (the number of Saturdays). In the first box on the second row write $10.48. This is how much Claire spends on 1 Saturday. In each of the next boxes on the second row, write the amount Claire spends by adding $10.48 to each previous amount. The answer in box 7 is the total amount Claire spent after 7 Saturdays. C. Guess and test. Guess that Claire spent $75. Multiply $3.99 by 2. Then add that answer to $2.50. Divide this answer into 75. If the answer is not 7, then revise your guess, and test your answer again.

OpenStudy (anonymous):

@phi

OpenStudy (phi):

Probably the table. But they're all lousy. You would say the mags are about 4 each 8+2.50= 10.50 * 7= 70 + 3.50 = 73.50 if you want to be exact, you overestimated by 2 cents per week or 14 cents so the exact answer is 73.36 but the first way you can do in your head (with practice)

OpenStudy (anonymous):

Okay, Thanks so much!!!!

OpenStudy (anonymous):

@maltitaly